40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Mar 13 03:12:47 2013, Manasa, Jenne, Update, Locking, TWO arms TWO colors 
    Reply  Wed Mar 13 08:34:33 2013, yuta, Update, Locking, TWO arms TWO colors 
       Reply  Wed Mar 13 16:29:25 2013, Manasa, Update, Locking, TWO arms TWO colors 
    Reply  Wed Mar 13 10:26:58 2013, Manasa, Update, Locking, We're still good!! 
    Reply  Mon Mar 25 15:01:33 2013, Manasa, Update, Locking, TT1 drift 
       Reply  Mon Mar 25 15:24:41 2013, rana, Update, Locking, TT1 drift 
Message ID: 8282     Entry time: Wed Mar 13 03:12:47 2013     Reply to this: 8283   8284   8336
Author: Manasa, Jenne 
Type: Update 
Category: Locking 
Subject: TWO arms TWO colors 

[Jenne, Manasa]

2 colors 2 arms realized!

1. Spot centering:

We spot centered the IR in both arms.
- Use TT1 and TT2 to center in Y arm (I visually center the spots on the ITM and ETM and then use TTs iteratively)
- Use BS-ETM to center in X arm

Spot positions after centering
               X arm            Y arm
         itmx    etmx        itmy    etmy
pitch    -0.86    0.37        1.51    0.05
yaw      0.01    -0.1        0.08    0.10


2. TT1 drifting in pitch (Bistable)
During the arm alignment routine for spot centering, we observed that TRY dropped (from TRY = 0.9 until the arm lost lock) every 40minutes or so. The arm was relocked by moving TT1 in pitch. The (locking - unlocking due to drift - relocking) cycle was monitored and we observed that it was bistable i.e. if TT1 was moved up in pitch (0.2 on the slider) to relock for the first time ; the next time it lost lock, TT1 had to be moved down by nearly the same distance to relock the arm.
Moving TT2 or the testmasses did not help with relocking the arms; so TT1 seems to be the one causing all th trouble atleast for today.

3. ALS - green alignment

We then moved on to Ygreen.  We used the out of vac steering mirrors to center the beam on the 2 irises that are in place on the table, which was a good starting place.  After doing that, and tweaking a small amount to overlap the incident and reflected beams on the green steering mirrors, we saw some mode lock.  We adjusted the end table steering mirrors until the Ygreen locked on TEM00.  We then followed Rana's suggestion of locking the IR to keep the cavity rigid while we optimized the green transmission.  Yuta, while adjusting ITMY and ETMY (rather than the out of vac mirrors) had been able to achieve a green transmission for the Yarm of ~2700 counts using the GTRX DC PD that's on the table. We were only able to get ~2200, with brief flashes up to 2500.

After that, we moved on to the X arm.  Since there are no irises on the table, we used the shutter as a reference, and the ETM optic itself.  Jenne looked through the viewport at the back of the ETM, while Manasa steered mirrors such that we were on the center of the ETM and the shutter.  After some tweaking, we saw some higher order modes lock.  We had a very hard time getting TEM00 to stay locked for more than ~1 second, even if the IR beam was locked.  It looks like we need to translate the beam up in pitch.  The leakage of the locked cavity mode is not overlapped with the incident beam or the promptly reflected beam.  This indicates that we're pretty far from optimally aligned.  Manasa was able to get up to ~2000 counts using the same GTRX PD though (with the Ygreen shutter closed, to avoid confusion).  Tomorrow we will get the Xarm resonating green in the 00 mode.

We need to do a little cleanup on the PSL green setup.  Yuta installed a shutter (I forget which unused one he took, but it was already connected to the computers.), so we can use it to block the PSL green beam.  The idea here is to use the 4th port of the combining beam splitters that are just before each beat PD, and place a PD and camera for each arm.  We already have 2 PDs on the table connected to channels, and one camera, so we're almost there. Jenne will work on this tommorrow during the day, so that we can try to get some beat signals and do some handoffs in the evening.

ELOG V3.1.3-